OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). a = 3b - 4 a + b =16

OpenStudy (anonymous):

a = 11, b=5 3b-4+b = 16 4b-4=16 4b=20 b=5 a=3(5)-4

OpenStudy (anonymous):

frm first eqn a = 3b-4 substitute in second eqn 3b - 4 + b = 16 4b = 16+4 b = 20/4 = 5

OpenStudy (anonymous):

a = 3b - 4 a + b =16 3b-4+b=16 4b=20 b=5 a=3(5)-4 a=15-4=11

OpenStudy (anonymous):

substitute b = 5 in second eqn a + 5 = 16 a = 16-5 a = 11

OpenStudy (anonymous):

i need the answer? please

OpenStudy (anonymous):

helpppppp

OpenStudy (anonymous):

your asking for it in (x,y) form but you give us a and b. is a=x and b=y? if so then (11,5) if not then we need to know more

OpenStudy (anonymous):

(a,b) = (11,5)